OpenStudy (anonymous):

: If you travel at the speed of 45 mph, what will this speed be in terms of miles per minute?

OpenStudy (anonymous):

\[\left( \frac{ 45 \text{miles} }{ \text{hour} } \right) \times \left( \frac{ 1 \text{hour} }{ 60 \text{minutes} } \right) \implies 0.75\text{mpm}\]